# Yu Wang  Juhyung Ha  David J. Crandall

### v2

@@ Abstract
Active speaker detection (ASD) in egocentric videos presents unique challenges due to unstable viewpoints, motion blur, and off-screen speech sources—conditions under which traditional visual-centric methods degrade significantly. We introduce PAIR-Net (Pretrained Audio–visual Integration with Regularization Network), an effective model that integrates a partially frozen Whisper audio encoder with a fine-tuned AV-HuBERT visual backbone to robustly fuse cross-modal cues. To counteract modality imbalance, we introduce an inter-modal alignment loss that synchronizes audio and visual representations, enabling more consistent convergence across modalities. Without relying on multi-speaker context or ideal frontal views, PAIR-Net achieves state-of-the-art performance on the Ego4D ASD benchmark with 76.6% mAP, surpassing LoCoNet and STHG by 8.2% and 12.9% mAP, respectively. Our results highlight the value of pretrained audio priors and alignment-based fusion for robust ASD under real-world egocentric conditions.
@@ 1 Introduction
Active speaker detection (ASD) is a fundamental task in multimodal understanding, supporting applications such as speaker diarization [14, 2], speech understanding [1], and human–robot interaction [18]. While prior methods [19, 8] perform well on third-person, high-quality footage with clearly visible faces, significant challenges remain in unconstrained settings. Specifically, egocentric videos from head-mounted cameras often suffer from motion blur, oblique angles, and low-resolution faces, while speech may originate from off-screen speakers, complicating attribution.
To address these challenges, recent methods such as LoCoNet [21] and STHG [11] have aimed to improve visual modeling by incorporating multi-speaker contextual reasoning, better accommodating the complexities of egocentric scenarios. While these approaches enhance visual robustness, they often underutilize the rich information embedded in the audio modality, limiting the full exploitation of cross-modal complementarities.
@@
We observe that the model often over-relies on visual cues while underutilizing audio, a phenomenon we term modality imbalance. To address this, we introduce an inter-modal alignment loss—motivated by LFM [23]—that enforces agreement between temporally aligned audio and visual features. This regularization balances convergence dynamics across modalities and, when combined with the standard classification loss, enhances overall performance. Notably, PAIR-Net achieves strong results using input from a single visible speaker, demonstrating both efficiency and generalization. On the Ego4D ASD benchmark [5], our method attains 76.6% mAP, surpassing prior state-of-the-art models LoCoNet [21] and STHG [11] by 8.2% and 12.9%, respectively.
@@ 3 Methodology
Figure 1: An overview of the proposed PAIR-Net. Modality-specific encoders process audio and visual inputs, with the early layers of the audio encoder frozen during training. The video encoder is followed by a trainable projector, while the audio encoder includes a temporal max pooling layer to align sequence lengths. A fully connected layer predicts the ASD label using the fused representation. Two additional modality-specific classifiers generate individual predictions, which are used to compute the inter-modal alignment loss.
We propose PAIR-Net, an audio-visual active speaker detection (ASD) framework that integrates powerful pretrained encoders with a simple yet effective fusion mechanism and contrastive supervision to mitigate modality imbalance.
3.1 Problem Formulation
Given a sequence of video frames $I_{v}\in\mathbb{R}^{T_{v}\times H\times W}$ containing cropped face regions and the corresponding log-Mel spectrogram $I_{a}\in\mathbb{R}^{T_{a}\times 80}$ derived from the raw audio, the goal is to predict frame-wise speaker activity.
@@ 3.2 Model Architecture
The PAIR-Net architecture is illustrated in Figure 1. We adopt the AV-HuBERT video encoder [17] and the Whisper [15] audio encoder for robust multimodal feature extraction. The video encoder is based on the AV-HuBERT visual branch [17], which consists of a ResNet backbone [6] for low-level feature extraction and 24 Transformer blocks for capturing high-level spatiotemporal features. For audio, we use the Whisper encoder [15], composed of two convolutional layers followed by 24 Transformer blocks, yielding temporally rich representations.
@@ 3.3 Modality Alignment
$\mathcal{L}_{\text{MM}}=\frac{1}{2}\text{KL}(y_{v}\|\tilde{y})+\frac{1}{2}\text{KL}(y_{a}\|\tilde{y}).$ (8)
@@ 4 Experiments
We evaluate our method on the Ego4D Audio-Visual Benchmark [5], focusing on Active Speaker Detection (ASD) task.

### v3

@@ Abstract
Active speaker detection (ASD) in egocentric videos presents unique challenges due to unstable viewpoints, motion blur, and off-screen speech sources—conditions under which traditional visual-centric methods degrade significantly. We introduce PAIR-Net (Pretrained Audio–visual Integration with Regularization Network), a model that integrates a partially frozen Whisper audio encoder with a fine-tuned AV-HuBERT visual backbone to robustly fuse cross-modal cues. To counteract modality imbalance, we introduce an inter-modal alignment loss that synchronizes audio and visual representations, enabling more consistent convergence across modalities. Without relying on multi-speaker context or ideal frontal views, PAIR-Net achieves state-of-the-art performance on the Ego4D ASD benchmark with 76.6% mAP, surpassing LoCoNet and STHG by 8.2% and 12.9% mAP, respectively. Our results highlight the value of pretrained audio priors and alignment-based fusion for robust ASD under real-world egocentric conditions.
@@ 1 Introduction
Active speaker detection (ASD) is a fundamental task in multimodal understanding, supporting applications such as speaker diarization [14, 2], speech understanding [1], and human–robot interaction [18]. While prior methods [19, 8] perform well on third-person, high-quality footage with clearly visible faces, significant challenges remain in unconstrained settings. Specifically, egocentric videos from head-mounted cameras often suffer from motion blur, oblique angles, and low-resolution faces, while speech may originate from off-screen speakers.
To address these challenges, recent methods such as LoCoNet [21] and STHG [11] have aimed to improve visual modeling by incorporating multi-speaker contextual reasoning. While these approaches enhance visual robustness, they often underutilize the rich information embedded in the audio modality, limiting the full exploitation of cross-modal complementarities.
@@
We observe that the model often over-relies on visual cues while underutilizing audio, a phenomenon we term modality imbalance. To address this, we introduce an inter-modal alignment loss—motivated by LFM [23]—that enforces agreement between temporally aligned audio and visual features. This regularization balances across modalities and, when combined with the standard classification loss, enhances overall performance. PAIR-Net achieves strong results using input from a single visible speaker, demonstrating both efficiency and generalization. On the Ego4D ASD benchmark [5], we attain 76.6% mAP, surpassing prior state-of-the-art models LoCoNet [21] and STHG [11] by 8.2% and 12.9%, respectively.
@@ 3 Methodology
Figure 1: An overview of PAIR-Net. Modality-specific encoders process audio and visual inputs, with the early layers of the audio encoder frozen during training. The video encoder is followed by a trainable projector, while the audio encoder includes a temporal max pooling layer to align sequence lengths. A fully connected layer predicts the ASD label using the fused representation. Two additional modality-specific classifiers generate individual predictions, which are used to compute the inter-modal alignment loss.
We propose PAIR-Net, an audio-visual active speaker detection (ASD) framework that combines powerful pretrained encoders with a simple yet effective fusion mechanism and contrastive supervision to mitigate modality imbalance. Given a sequence of video frames $I_{v}\in\mathbb{R}^{T_{v}\times H\times W}$ containing cropped face regions and the corresponding log-Mel spectrogram $I_{a}\in\mathbb{R}^{T_{a}\times 80}$ derived from the raw audio, the goal is to predict frame-wise speaker activity.
@@ 3.1 Model Architecture
The PAIR-Net architecture is illustrated in Figure 1. We use the AV-HuBERT video encoder [17] and the Whisper [15] audio encoder for robust multimodal feature extraction. The video encoder is based on the AV-HuBERT visual branch [17], which consists of a ResNet backbone [6] for low-level feature extraction and 24 Transformer blocks for capturing high-level spatiotemporal features. For audio, we use the Whisper encoder [15], composed of two convolutional layers followed by 24 Transformer blocks, yielding temporally rich representations.
@@ 3.2 Modality Alignment
[…]
@@ 4 Experiments
We evaluate on the Ego4D Audio-Visual Benchmark [5], focusing on Active Speaker Detection (ASD).
@@ 5 Conclusion
6 Acknowledgments
This work was supported in part by the National Science Foundation under award DRL-2112635 to the AI Institute for Engaged Learning. Any opinions, findings, conclusions, or recommendations are those of the author(s) and do not necessarily reflect those of the NSF.